OpenStudy (anonymous):

wat is 2/3moltiplied by 4 minus 4?????? help me pleeeeze

OpenStudy (anonymous):

come on

OpenStudy (anonymous):

suck my wingspan polpak!

OpenStudy (anonymous):

pol pak flutter you motherflutterer

OpenStudy (anonymous):

immature jonblabla! (2/3)*4-4=8/3-4=8/3-12/3=-4/3

OpenStudy (anonymous):

polpak suck my wingspan

OpenStudy (anonymous):

fluttering feather retricenerd

OpenStudy (anonymous):

I can only assume that you're gay by your instant requests that I provide you with oral sex. I'm not gay though, so thanks anyway.

OpenStudy (anonymous):

Incessant rather.

OpenStudy (anonymous):

lol i was talking about ur mom owletg

OpenStudy (anonymous):

she can give a mean retriceblowjob

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

My mother is quite elderly. I didn't realize you were into geriatrics.

OpenStudy (anonymous):

hey pussy hella juicy

OpenStudy (anonymous):

she takes off those dentures and get to sucking mean strokes it hard too

OpenStudy (anonymous):

shut the flutter up how bout yall just go jack off to eachother

OpenStudy (anonymous):

Jonathan, perhaps you should go ask your mother why she didn't raise you properly instead of wasting your time trying to pick up on underage girls in an online forum for math homework..

OpenStudy (anonymous):

conner, there's the answer somewhere in the middle of their conversation lol

OpenStudy (anonymous):

wow im 14

OpenStudy (anonymous):

lets have sex and beat eachothers meat

OpenStudy (anonymous):

do your math right wingspansucking idiot

OpenStudy (anonymous):

tell ur mom i owe her 2 dollars

OpenStudy (anonymous):

get off my fluttering math thing u feather retricenerds

OpenStudy (anonymous):

and tell her too give me back my pants

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

can i have one pp touch

OpenStudy (anonymous):

im just here too annoy nerds like u and polapak connor! owletgots

OpenStudy (anonymous):

Connor, you got your answer already.. I don't know why you still care. I'm only here because otherwise he'll follow me somewhere else and ruin someone elses problem.

OpenStudy (anonymous):

stupid mother flutterers i will beet all your asses if u dont shut the hell up

OpenStudy (anonymous):

owletgot im so scared !!!

OpenStudy (anonymous):

your mom was scared when i ripped her anal

OpenStudy (anonymous):

both of your mom sucked my fat retricewingspan

OpenStudy (anonymous):

and licked my retrice

OpenStudy (anonymous):

she gave me a blumpkin

OpenStudy (anonymous):

im not a nerd i got in a fight and i got suspended so i dont kno how to do this pellet

OpenStudy (anonymous):

and tell her to clean her blue waffle and cut her roast beef curtains

OpenStudy (anonymous):

i got expelled conner :). ur answer here (2/3)*4-4=8/3-4=8/3-12/3=-4/3

OpenStudy (anonymous):

Connor. which part of ttt's explaination didn't you understand? Also I don't know that calling people who are trying to help you 'nerds' is terribly polite

OpenStudy (anonymous):

flutter off you whores